given that cosA=1/3 where A is an acute angle, the exact value of cos A/2 equals
Answer by lwsshak3(11628)

given that cosA=1/3 where A is an acute angle, the exact value of cos A/2 equals
***
use half-angle formula for cos:
cos(A/2)=√[(1+cosA)/2]=√(1+(1/3))/2)=√((4/3)/2)=√(4/6)=√(2/3)
Check with calculator:
cosA=1/3
A≈70.5288º
A/2≈35.2644º
cos(A/2)≈cos35.2644º≈0.8165
√(2/3)≈0.8165